# Improving the efficiency of finite-time memory erasure with potential barrier shaping

### v4

@@ III.1 Success rate
First, we analyze the efficiency of the erasure mechanism concerning erasure rate or the success rate of the process, which is the ratio of the successful erasure cycles and the total number of cycles for the ensemble. To proceed, we have simulated the overdamped Langevin equation (6) using an improved Euler algorithm [press1992numerical]. Gaussian white noise $\xi(t)$ is generated using the Box–Muller algorithm. First, we consider the case of starting the erasure protocol with equilibrium memory. In these simulations, at the start of the time, we place all the particles of the ensemble at position $x=0$, i.e., at the barrier top, and let the system get thermalized with the heat bath at temperature $T$ and arrive at their steady-state distribution according to the structure of the underlying potential. As we consider a range of asymmetric forms of the potential, we get unique steady-state distributions after the initial thermalization period, as discussed above. We have examined the attainment of this steady-state distribution for all cases studied. To express the scenario in terms of the representation of the memory states, we can say that both the memory states get occupied after the thermalization period, following a steady-state distribution specific to the characteristic asymmetry of the governing potential. For a symmetric potential, this leads to an equal population of both memory states, “0” and “1” just before the start of the erasure protocol.
Memory erasure occurs when the particle makes a passage from one state to the other. When the barrier height is significantly greater than the thermal energy $k_{B}T$, the particle remains confined in one of the potential wells, i.e., at one of the memory states, and is unable to transition between them. To make this passage happen, we use a process where we lower the barrier height and apply a tilting force at the same time after the particle has had time to relax. [D_and_Lutz, Jun_prl_2014, PhysRevResearch_Jun]. In our applied protocol, we aim for the particles to reside in the right well at the end of the process. The tilting force is applied accordingly in the appropriate direction. Therefore, following this mechanism, the goal is to erase the memory state 0, so that ultimately, the memory state $1$ is occupied, starting from an initial steady-state distribution where both the memory states exist.
@@
To quantify the above understanding, we plot the erasure rate or the success rate of the erasure process, defined as the fraction of particles that occupy the target state, i.e., the memory state 1, represented by the right well of the potential, by the end of the protocol, as a function of the driving amplitude $(A)$. The results are obtained from an ensemble of $10^{5}$ particles. We observe that in the presence of greater asymmetry, a significantly smaller external driving force is sufficient to achieve a 100% erasure rate, in contrast to the symmetric case, as shown in Fig. 5(a)-(b) for two different initial conditions starting with the equilibrium and nonequilibrium memory states, respectively. In both of the situations, we get similar observations.
@@
Figure 5: The erasure performance as a function of the driving amplitude of the tilting force $A$ is illustrated. Panels (a) and (b) show how the erasure rate changes with AA for different asymmetry parameters $c$. In both panels, barrier height $h=1$, noise strength $D=0.15$, the barrier-lowering depth is $Z=0.5$, $\omega=0.01$, $\tau_{1}=298$, and $\tau_{2}=16$. In panel (a), the results are for the erasure mechanisms when the particles are initially positioned at $x=0$ at the start of the simulations. Then they are allowed to get thermalized with the heat bath for the duration $t_{0}=314$, whereas for the data in panel (b), the particles have an initial 50-50 distribution without thermalization. For both cases, the ramp-up and ramp-down times for the tilting force are $\tau_{1}=298$ and $\tau_{2}=16$, respectively.
@@
Next, we have shown how the erasure rate changes with another important quantity influencing the dynamics, the noise strength $D$ in Fig. 6(a), while keeping the external force constant at $A=1.50$, for different extent of asymmetries in the intrinsic potential. We detect that as the noise strength increases, the erasure rate decreases, reaching a saturation point in the higher range of $D$. This behavior at higher noise strength can be attributed to spontaneous thermal fluctuations that increasingly drive the particles to the wrong well, or the memory state, thereby reducing the effectiveness of the erasure.
@@
We understand from the previous results of the variation of the erasure rate or the success rate as a function of the tilting force amplitude $A$ and the noise strength $D$, for different values of the asymmetry parameter $c$, that the erasure rate is highly sensitive to the degree of asymmetry present in the potential. Therefore, we perform a more direct study to observe the functional dependence of the erasure rate on the asymmetry parameter $c$. This explicit variation of the erasure rate against $c$ is plotted in Fig. 6(b), where we show how the erasure rate changes with the asymmetry parameter $c$ for different levels of the external driving force amplitude $A$ for a fixed value of noise strength $D$. For lower values of $A$, we observe a higher erasure rate in the strongly asymmetric regime, but as the system approaches symmetry ($c\to 1$), the erasure rate drops significantly. However, for sufficiently high external forces (e.g., $A=1.30$), the erasure rate remains high (around 95%) even in the symmetric case ($c=1$) and does not vary substantially as the asymmetry is reduced. This observation suggests that the effect of asymmetry becomes more prominent when the amplitude of the tilting force is less. This further justifies the fact that the asymmetry factor in the potential makes the transition favorable, which was less probable due to the impact of the erasure protocol alone.
@@
Figure 6: Panel (a) shows the erasure rate as a function of the noise strength $D$, for different asymmetry parameter values with fixed barrier height $h=1$, tilting force amplitude $A=1.50$, protocol shape parameter $Z=0.5$ and $\omega=0.01$. Panel (b) presents the erasure rate as a function of the asymmetry parameter $c$, for fixed values of the barrier height $h=1$, noise strength $D=0.15$, protocol shape parameter $Z=0.5$, and $\omega=0.01$. For the results in both panels, the particles are placed at $x=0$ at the start of the simulations and are allowed to get thermalized with the heat bath. The thermalization time is $t_{0}=314$, and the ramp-up and ramp-down durations of the tilting force are $\tau_{1}=298$ and $\tau_{2}=16$, respectively.
@@
Erasing from memory-state “1” to memory-state “0”: To further confirm this explanation, we also performed the reverse erasing mechanism, where finally the left well or the memory state 0 is expected to be occupied. The variation of the success rate or the erasure rate as a function of the tilting amplitude $A$ is plotted in Figs. 8(a)-(b), for the systems with initial equilibrium and 50:50 distributions of memory states, respectively. In this case of the reverse protocol, it is observed that a larger tilting force is required to perform the erasure when the system has asymmetry. This again supports our statement concerning the greater preference for the occupancy of the higher entropy state.
In all of the above-mentioned investigations, we verify the findings based on the initial equilibrium and nonequilibrium conditions. Our studies reveal that the observations are similar for both cases. These facts point towards the generality of the effect of the asymmetry on the erasure rate, irrespective of starting the protocol with equilibrium or nonequilibrium memory states.
@@
Figure 8: The erasure rate as a function of the tilting force amplitude $A$ for the reverse mechanism for different asymmetry parameters $c$. The reverse protocol refers to the scenario where, after the successful erasure, the particle arrives at the left well, i.e., at the memory state 0. The parameters are fixed at $h=1$, $D=0.15$, $z=0.5$, and $\omega=0.01$. The results have been presented for the protocols with initial (a) equilibrium and (b) 50:50 distribution of memory states. For case (a), the system undergoes thermalization for a duration of $t_{0}=314$. No thermalization period is allowed for case (b). In both situations, the protocol phase consists of the ramp-up and ramp-down durations of $\tau_{1}=298$ and $\tau_{2}=16$, respectively, for the tilting force.
@@ III.2 Heat and work
Next, we focus on the essential thermodynamic quantities required to describe and interpret the efficiency of the erasure process from the perspective of energetics. These quantities are mainly the work done $W$ and the evolved heat $Q$ associated with the erasure mechanism. We proceed to estimate these observables by employing the fundamental relations of thermodynamics. We implement the specific characteristics of the system and dynamics for our present study in these relations while doing the measurements. The change in the internal energy or the total potential energy is another fundamental energy function that contributes to the development of a detailed description of the energetic aspects of the erasure process. It is also necessary to depict the connection between the work and heat involvement for the protocol. The change in the total potential energy of the Brownian particle during the erasure process is defined as [Sekimoto_1998, udo_2012],
@@
where U(x,t)=V(x,t)xAf(t)U(x,t)=V(x,t)-xAf(t) represents the intrinsic potential combined with the potential appearing due to the external driving force. $V(x,t)$ can have an asymmetric or a symmetric form depending on the value of the parameter $c$ present in its expression, as discussed before. The total work $W$ performed on the system during the erasure process can be calculated considering the change in the internal energy as [Moupriya_PhysRevResearch, udo_2008],
@@
We estimate $\Delta U$ and $W$ for each trajectory using the numerical solution of $x$ according to the Langevin dynamics (Eq. (1)). These two thermodynamic functions are calculated by employing Eq. (7) and Eq. (9), respectively. According to the first law of thermodynamics, the heat dissipated into the thermal bath $Q=\Delta U-W$. Therefore, $Q$ can be measured from the solutions of Eq. (7) and Eq. (9). As we are dealing with small systems subject to thermal fluctuations, all these quantities, $\Delta U$, $W$, and $Q$, are stochastic variables for a single trajectory. Their ensemble-averaged values, $\langle\Delta U\rangle$, $\langle W\rangle$, and $\langle Q\rangle$ are considered the reliable statistical measures of the quantities and are referred to for further discussions. Also, the first law of thermodynamics takes the form as $\langle\Delta U\rangle=\langle Q\rangle+\langle W\rangle$ for small systems [peliti2021stochastic, H_Kantz, toyabe2010experimental, hasegawa2010generalization, Ueda_2014_pre, Supriya_prr_2024, RevModPhys.97.015002]. Since we have a periodic driving force that satisfies $f(t_{0}+t_{f})=f(t_{0})=0$, the net contribution of the external drive vanishes over the duration of a period. During a complete erasure cycle, the average change in potential energy remains unaltered, that is, $\langle\Delta U\rangle=0$. Therefore, the mean work performed on the system equals the heat dissipated into the bath: $|\langle W\rangle|=|\langle Q\rangle|$.
@@
Figure 9: Work and heat distributions for different asymmetry parameters: (a) and (b) correspond to $c=1$ and $A=1.30$; (c) and (d) correspond to $c=0.74$ and $A=0.90$; (e) and (f) correspond to $c=0.56$ and $A=0.75$. The barrier height of the potential is $h=1$ and the noise strength $D=0.15$. The parameters of the barrier lowering force are $\omega=0.01$, $Z=0.5$, The ramp-up and ramp-down durations of the tilting force are $\tau_{1}=298$, and $\tau_{2}=16$, respectively. The results in panels (a), (c), and (e) are for initial equilibrium distributions of the memory states; in panels (b), (d), and (f), we represent the results when particles are initially distributed equally between both wells (50-50 distribution) without thermalization at the start of the erasure protocol. The Landauer bound $D\ln 2$ is shown with the solid straight line. $\langle W\rangle$ and $\langle Q\rangle$ are indicated with dashed lines.
Figure 10: Mean work $\langle W\rangle$ and heat $\langle Q\rangle$ versus protocol duration $\tau$ for a fixed potential height $h=1$, noise strength $D=0.15$, barrier lowering parameters $\omega=\pi/\tau$, and $z=0.5$. The system is first thermalized for a duration of $t_{0}=314$, followed by a protocol phase with a ramp-up and ramp-down time ratio of $\tau_{1}:\tau_{2}=19:1$ for the tilting force. The results have been presented for four asymmetry parameter values $c=1,0.56,0.34,\textrm{and}\;\ 0.2$. The corresponding $A_{\textrm{opt}}$ has been chosen as the amplitude of the tilting force for each case.
@@
Next, we analyzed the mean work $\langle W\rangle$ and heat $\langle Q\rangle$ as a function of protocol duration $\tau$ under the influence of both external driving force and barrier modulation, with $\tau_{1}:\tau_{2}=19:1$. While many experimental studies  [berut_2013] have neglected the barrier modulation contribution, we explicitly account for both functions in our analysis. It is understood that the average value of the work and heat decreases with the increasing value of $\tau$ as we approach the quasi-reversibility  [berut_lutz2012nature, moupriya_2014]. This is observed for erasure mechanisms in different cases with the symmetric and asymmetric potentials. Also, it is known that in the very long-time limit, these average quantities advance towards the Landauer bound for the erasure process in the symmetric case [berut_lutz2012nature, Moupriya_PhysRevE_2014, moupriya_2014]. This is what has again been retrieved in our present analysis of erasure for the $c=1$ case. Interestingly, for the asymmetric cases, we observe a violation of the Landauer bound. The results have been shown in Fig. 10. We compare the findings for the equilibrium and nonequilibrium initial memory states. The observations for the nonequilibrium initial memory condition have been presented in Fig. 11. We detect that the basic interpretations remain the same for both circumstances. Sagawa and Ueda [Sagawa_2009] previously demonstrated that Landauer’s principle can be violated when the initial state is in nonequilibrium and the system exhibits asymmetry. This theoretical prediction was subsequently verified experimentally by Ciampini et al. [ciampini2021experimental]. However, in our present study, we also allow the system to reach equilibrium before initiating the protocol, suggesting that asymmetry alone can lead to violations of the Landauer bound even under equilibrium initial conditions. Moreover, we explore in detail how the extent of asymmetry can affect the average values of work and heat associated with the finite-time erasure processes. To develop a more systematic understanding of this fact, we scrutinize the variation of $\langle W\rangle$ and $\langle Q\rangle$ with the asymmetry parameter $c$ for a fixed value of the amplitude of the erasure force $A$. We find that for higher asymmetry, we need to do much less work compared to the symmetric case. The results for three fixed values of $A$ are shown in Fig. 12. The same has been examined for the erasure protocols with nonequilibrium initial conditions of the memory states. The results have been illustrated in Fig. 13. From the comparisons of the data in Fig. 12 and Fig. 13, we detect a monotonic rise of $\langle W\rangle\$and $\langle Q\rangle\$as a function of $c$ in both cases. However, the rate of the rise is steeper in the case of the equilibrium initial memory condition.

### v5

@@ III.1 Success rate
First, we analyze the efficiency of the erasure mechanism concerning the success rate of the process, which is the ratio of the successful erasure cycles and the total number of cycles for the ensemble. To proceed, we have simulated the overdamped Langevin equation (6) using an improved Euler algorithm [58]. Gaussian white noise $\xi(t)$ is generated using the Box–Muller algorithm. First, we consider the case of starting the erasure protocol with equilibrium memory. In these simulations, at the start of the time, we place all the particles of the ensemble at position $x=0$, i.e., at the barrier top, and let the system get thermalized with the heat bath at temperature $T$ and arrive at their steady-state distribution according to the structure of the underlying potential. As we consider a range of asymmetric forms of the potential, we get unique steady-state distributions in two wells after the initial thermalization period, as discussed above. We have examined the attainment of this steady-state distribution for all cases studied.To express the scenario in terms of the representation of the memory states, we can say that both the memory states get occupied after the thermalization period, following a steady-state distribution specific to the characteristic asymmetry of the governing potential. For a symmetric potential, this leads to an equal population of both memory states, “0” and “1” just before the start of the erasure protocol.
In the case of the erasure of binary memories, one of the memory states vanishes at the end of the process, while initially, both memory states are present. This is an essential step of the irreversible logic operations, in which many (at least two) input memory states are mapped onto one output memory state, thereby erasing information of the starting condition. In a physical setup, this mechanism is modeled in a bistable potential where the states of an ensemble of Brownian particles represent the binary memory 0 and $1$. The erasure force is applied to the particles, which are initially distributed in both wells, so as to drive all the particles in the target well or to the desired memory state at the end of the protocol. As the particles in the ensemble are indistinguishable, the information about the starting position for each particle is lost during the erasure mechanism. When the barrier height of the potential is significantly greater than the thermal energy $k_{B}T$, the particle remains confined in one of the potential wells, i.e., at one of the memory states, and is unable to transition between them. To make the passage happen to the target well, if the particle is initially at the other well, we use a process where we lower the barrier height and apply a tilting force at the same time after the particle has had time to relax. [15, 60, 61]. In our applied protocol, we aim for the particles to reside in the right well at the end of the process. The tilting force is applied accordingly in the appropriate direction. Therefore, following this mechanism, the goal is to erase the memory state 0, so that ultimately, the memory state $1$ is occupied, starting from an initial steady-state distribution where both the memory states exist.
@@
To quantify the above understanding, we plot the success rate of the erasure process, defined as the fraction of particles that occupy the target state, i.e., the memory state 1, represented by the right well of the potential, by the end of the protocol, as a function of the driving amplitude $(A)$. The results are obtained from an ensemble of $10^{5}$ particles. We observe that in the presence of greater asymmetry, a significantly smaller external driving force is sufficient to achieve a 100% success rate, in contrast to the symmetric case, as shown in Fig. 6(a)-(b) for two different initial conditions starting with the equilibrium and nonequilibrium memory states, respectively. In both of the situations, we get similar observations.
@@
Figure 6: The erasure performance as a function of the driving amplitude of the tilting force $A$ is illustrated. Panels (a) and (b) show how the success rate changes with AA for different asymmetry parameters $c$. In both panels, barrier height $h=1$, noise strength $D=0.15$, the barrier-lowering depth is $Z=0.5$, $\omega=0.01$, $\tau_{1}=298$, and $\tau_{2}=16$. In panel (a), the results are for the erasure mechanisms when the particles are initially positioned at $x=0$ at the start of the simulations. Then they are allowed to get thermalized with the heat bath for the duration $t_{0}=314$, whereas for the data in panel (b), the particles have an initial 50-50 distribution without thermalization. For both cases, the ramp-up and ramp-down times for the tilting force are $\tau_{1}=298$ and $\tau_{2}=16$, respectively.
@@
Next, we have shown how the success rate changes with another important quantity influencing the dynamics, the noise strength $D$ in Fig. 7(a), while keeping the external force constant at $A=1.50$, for different extent of asymmetries in the intrinsic potential. We detect that as the noise strength increases, the success rate decreases, reaching a saturation point in the higher range of $D$. This behavior at higher noise strength can be attributed to spontaneous thermal fluctuations that increasingly drive the particles to the wrong well, or the memory state, thereby reducing the effectiveness of the erasure.
@@
We understand from the previous results of the variation of the success rate or the success rate as a function of the tilting force amplitude $A$ and the noise strength $D$, for different values of the asymmetry parameter $c$, that the success rate is highly sensitive to the degree of asymmetry present in the potential. Therefore, we perform a more direct study to observe the functional dependence of the success rate on the asymmetry parameter $c$. This explicit variation of the success rate against $c$ is plotted in Fig. 7(b), where we show how the success rate changes with the asymmetry parameter $c$ for different levels of the external driving force amplitude $A$ for a fixed value of noise strength $D$. For lower values of $A$, we observe a higher success rate in the strongly asymmetric regime, but as the system approaches symmetry ($c\to 1$), the success rate drops significantly. However, for sufficiently high external forces (e.g., $A=1.30$), the success rate remains high (around 95%) even in the symmetric case ($c=1$) and does not vary substantially as the asymmetry is reduced. This observation suggests that the effect of asymmetry becomes more prominent when the amplitude of the tilting force is less. This further justifies the fact that the asymmetry factor in the potential makes the transition favorable, which was less probable due to the impact of the erasure protocol alone.
@@
Figure 7: Panel (a) shows the success rate as a function of the noise strength $D$, for different asymmetry parameter values with fixed barrier height $h=1$, tilting force amplitude $A=1.50$, protocol shape parameter $Z=0.5$ and $\omega=0.01$. Panel (b) presents the success rate as a function of the asymmetry parameter $c$, for fixed values of the barrier height $h=1$, noise strength $D=0.15$, protocol shape parameter $Z=0.5$, and $\omega=0.01$. For the results in both panels, the particles are placed at $x=0$ at the start of the simulations and are allowed to get thermalized with the heat bath. The thermalization time is $t_{0}=314$, and the ramp-up and ramp-down durations of the tilting force are $\tau_{1}=298$ and $\tau_{2}=16$, respectively.
@@
Erasing from memory-state “1” to memory-state “0”: To further confirm this explanation, we also performed the reverse erasing mechanism, where finally the left well or the memory state 0 is expected to be occupied. The variation of the success rate or the success rate as a function of the tilting amplitude $A$ is plotted in Figs. 9(a)-(b), for the systems with initial equilibrium and 50:50 distributions of memory states, respectively. In this case of the reverse protocol, it is observed that a larger tilting force is required to perform the erasure when the system has asymmetry. This again supports our statement concerning the greater preference for the occupancy of the higher entropy state.
In all of the above-mentioned investigations, we verify the findings based on the initial equilibrium and nonequilibrium conditions. Our studies reveal that the observations are similar for both cases. These facts point towards the generality of the effect of the asymmetry on the success rate, irrespective of starting the protocol with equilibrium or nonequilibrium memory states.
@@
Figure 9: The success rate as a function of the tilting force amplitude $A$ for the reverse mechanism for different asymmetry parameters $c$. The reverse protocol refers to the scenario where, after the successful erasure, the particle arrives at the left well, i.e., at the memory state 0. The parameters are fixed at $h=1$, $D=0.15$, $z=0.5$, and $\omega=0.01$. The results have been presented for the protocols with initial (a) equilibrium and (b) 50:50 distribution of memory states. For case (a), the system undergoes thermalization for a duration of $t_{0}=314$. No thermalization period is allowed for case (b). In both situations, the protocol phase consists of the ramp-up and ramp-down durations of $\tau_{1}=298$ and $\tau_{2}=16$, respectively, for the tilting force.
@@ III.2 Heat and work
Next, we focus on the essential thermodynamic quantities required to describe and interpret the efficiency of the erasure process from the perspective of energetics. These quantities are mainly the work done $W$ and the evolved heat $Q$ associated with the erasure mechanism. We proceed to estimate these observables by employing the fundamental relations of thermodynamics. We implement the specific characteristics of the system and dynamics for our present study in these relations while doing the measurements. The change in the internal energy or the total potential energy is another fundamental energy function that contributes to the development of a detailed description of the energetic aspects of the erasure process. It is also necessary to depict the connection between the work and heat involvement for the protocol. The change in the total potential energy of the Brownian particle during the erasure process is defined as [62, 63],
@@
where U(x,t)=V(x,t)xAf(t)U(x,t)=V(x,t)-xAf(t) represents the intrinsic potential combined with the potential appearing due to the external driving force. $V(x,t)$ can have an asymmetric or a symmetric form depending on the value of the parameter $c$ present in its expression, as discussed before. The total work $W$ performed on the system during the erasure process can be calculated considering the change in the internal energy as [64, 65],
@@
We estimate $\Delta U$ and $W$ for each trajectory using the numerical solution of $x$ according to the Langevin dynamics (Eq. (1)). These two thermodynamic functions are calculated by employing Eq. (7) and Eq. (9), respectively. According to the first law of thermodynamics, the heat dissipated into the thermal bath $Q=\Delta U-W$. Therefore, $Q$ can be measured from the solutions of Eq. (7) and Eq. (9). As we are dealing with small systems subject to thermal fluctuations, all these quantities, $\Delta U$, $W$, and $Q$, are stochastic variables for a single trajectory. Their ensemble-averaged values, $\langle\Delta U\rangle$, $\langle W\rangle$, and $\langle Q\rangle$ are considered the reliable statistical measures of the quantities and are referred to for further discussions. Also, the first law of thermodynamics takes the form as $\langle\Delta U\rangle=\langle Q\rangle+\langle W\rangle$ for small systems [66, 67, 68, 69, 70, 71, 72]. Since we have a periodic driving force that satisfies $f(t_{0}+t_{f})=f(t_{0})=0$, the net contribution of the external drive vanishes over the duration of a period. During a complete erasure cycle, the average change in potential energy remains unaltered, that is, $\langle\Delta U\rangle=0$. Therefore, the mean work performed on the system equals the heat dissipated into the bath: $|\langle W\rangle|=|\langle Q\rangle|$.
@@
Figure 10: Work and heat distributions for different asymmetry parameters: (a) and (b) correspond to $c=1$ and $A=1.30$; (c) and (d) correspond to $c=0.74$ and $A=0.90$; (e) and (f) correspond to $c=0.56$ and $A=0.75$. The barrier height of the potential is $h=1$ and the noise strength $D=0.15$. The parameters of the barrier lowering force are $\omega=0.01$, $Z=0.5$, The ramp-up and ramp-down durations of the tilting force are $\tau_{1}=298$, and $\tau_{2}=16$, respectively. The results in panels (a), (c), and (e) are for initial equilibrium distributions of the memory states; in panels (b), (d), and (f), we represent the results when particles are initially distributed equally between both wells (50-50 distribution) without thermalization at the start of the erasure protocol. In the sub-figures, the Landauer bound $D\ln 2$ is shown with the dashed line. $\langle W\rangle$ and $\langle Q\rangle$ are indicated with solid straight line and $\Delta F_{\textrm{effective}}$ is presented with dotted dash line.
Figure 11: Mean work $\langle W\rangle$ and heat $\langle Q\rangle$ versus protocol duration $\tau$ for a fixed potential height $h=1$, noise strength $D=0.15$, barrier lowering parameters $\omega=\pi/\tau$, and $z=0.5$. The system is first thermalized for a duration of $t_{0}=314$, followed by a protocol phase with a ramp-up and ramp-down time ratio of $\tau_{1}:\tau_{2}=19:1$ for the tilting force. The results have been presented for four asymmetry parameter values $c=1,0.56,0.34,\textrm{and}\;\ 0.2$. The corresponding $A_{\textrm{opt}}$ has been chosen as the amplitude of the tilting force for each case. The estimated values of $\Delta F_{\textrm{effective}}$ with error bars, calculated in the given range of $\tau$, are plotted for all values of $c$, to compare with the corresponding $\langle W\rangle$ and $\langle Q\rangle$.
@@
Next, we analyzed the mean work $\langle W\rangle$ and heat $\langle Q\rangle$ as a function of protocol duration $\tau$ under the influence of both external driving force and barrier modulation, with $\tau_{1}:\tau_{2}=19:1$. While many experimental studies  [73] have neglected the barrier modulation contribution, we explicitly account for both functions in our analysis. It is understood that the average value of the work and heat decreases with the increasing value of $\tau$ as we approach the quasi-reversibility  [16, 74]. This is observed for erasure mechanisms in different cases with the symmetric and asymmetric potentials. Also, it is known that in the very long-time limit, these average quantities advance towards the Landauer bound for the erasure process in the symmetric case [16, 75, 74]. This is what has again been retrieved in our present analysis of erasure for the $c=1$ case. Interestingly, for the asymmetric cases, we observe a violation of the Landauer bound. The results have been shown in Fig. 11. We compare the findings for the equilibrium and nonequilibrium initial memory states. The observations for the nonequilibrium initial memory condition have been presented in Fig. 12. We detect that the basic interpretations remain the same for both circumstances. Sagawa and Ueda [52] previously demonstrated that Landauer’s principle can be violated when the initial state is in nonequilibrium and the system exhibits asymmetry. This theoretical prediction was subsequently verified experimentally by Ciampini et al. [49]. However, in our present study, we also allow the system to reach equilibrium before initiating the protocol, suggesting that asymmetry alone can lead to violations of the Landauer bound even under equilibrium initial conditions. Moreover, we explore in detail how the extent of asymmetry can affect the average values of work and heat associated with the finite-time erasure processes. To develop a more systematic understanding of this fact, we scrutinize the variation of $\langle W\rangle$ and $\langle Q\rangle$ with the asymmetry parameter $c$ for a fixed value of the amplitude of the erasure force $A$. We find that for higher asymmetry, we need to do much less work compared to the symmetric case. The results for three fixed values of $A$ are shown in Fig. 13. The same has been examined for the erasure protocols with nonequilibrium initial conditions of the memory states. The results have been illustrated in Fig. 14. From the comparisons of the data in Fig. 13 and Fig. 14, we detect a monotonic rise of $\langle W\rangle\$and $\langle Q\rangle\$as a function of $c$ in both cases. However, the rate of the rise is steeper in the case of the equilibrium initial memory condition.